# HPE-CogVLM: New Head Pose Grounding Task Exploration on Vision Language Model

### v1

@@ 3 HPE-CogVLM Framework
At this initial stage, the original grounding CogVLM undergoes pre-training on the CrowdHuman dataset. The function of this stage is to enhance the model’s understanding of the HPE task, as the CrowdHuman dataset offer a rich collection of human head images. Since the original CrowdHuman dataset does not provide ground truth (GT) annotations for HPE, we infer the weak HPE annotations by 6DRepNet. The output model from this stage is termed as weak label CogVLM as shown in Figure 2. For this stage, our primary goal is to warm up the model using weak label data, designed to provide the model with a comprehensive understanding of various human head orientations.
@@
Stage 2: Supervised Fine-tuning of the Weak Label CogVLM on Task-specific (HPE) Data
Following the pre-training, the model progresses to supervise fine-tuning solely using the task-specific HPE dataset. The task-specific HPE datasets contain less images and more accurate annotations than the weak label images. This stage concentrates on refining the weak label model’s HPE ability, aiming to maximize the precision of HPE task. The output model is referred as the HPE-oriented CogVLM as shown in Figure 2.
@@
Stage 3: Layer-based Merging between Original Grounding CogVLM and HPE-oriented CogVLM
@@
During this key stage, the original grounding CogVLM is merged with HPE-oriented CogVLM based on cosine similarity criteria. In our framework, the cosine similarity represents the averaged cosine similarity between layer parameter tensors along the last dimension. Cosine similarity is used to gauge the amount of information shared between layers. A high threshold of cosine similarity is set to ensure significant overlap in content. If the similarity falls below this threshold, we opt to completely retain the original knowledge. Otherwise, if the similarity exceeds the threshold, which indicates a substantial overlap in information due to the stringent criteria, we select the entire layer from the HPE-oriented CogVLM to guarantee the minimal risk of losing important existing knowledge.
The precedent methods merge models at the parameter-level by setting hyper parameters, or developing algorithms to discard and merge specific parameters ilharco2023editing ; yadav2024ties ; yu2023language ; akiba2024evolutionary . However, parameter-based merging models often blend output structures in our task, resulting in invalid answers. For example, when we query with HPE prompts, the parameter-based merging model may return a NLP response like “a person of head” or provide nonsensical responses like “[[999,231,123,389}”. More examples are detailed in Appendix A.1 Table 5. To overcome this issue, we adopt the “winner takes all” method to choose layers from either the original grounding CogVLM or the HPE-oriented CogVLM. The layer-based merging CogVLM is able to preserve the integrity of parameters specific to the expertise of each model by this approach. The merging criteria is detailed as below:
@@
  • When the cosine similarity between two layers from each model is less than the threshold (set at 0.95 in our experiments), we also select the layer from the original grounding CogVLM.
@@
Stage 4: Continual Fine-tuning of Layer-based Merging CogVLM on Mixture Data
After merging, the layer-based merging CogVLM undergoes an additional round of fine-tuning with both the task-specific HPE dataset and the rehearsal images. We pre-define the optimal rehearsal ratio from stage 1 by tuning the original grounding CogVLM with weak label images combined with different proportions of rehearsal images. Then the optimal rehearsal ratio is used in fine-tuning the merged model. Unlike the fine-tuning in stage 2, this phase involves only a brief period of fine-tuning, less than one epoch. The rationale for incorporating additional brief fine-tuning is that while layer merging maintains parameter integrity, it lacks the fine-tuned parameters necessary to enhance prediction accuracy. Continual fine-tuning is the best way to update weights than any other algorithms. In our approach, the merging model can be quickly fine-tuned to deliver accurate numerical predictions. The final output model of this stage is HPE-CogVLM as shown in Figure 2.
Stage 5: Evaluation of HPE-CogVLM on Test Data
@@
4 Experiments Setup
4.1 HPE Task Prompt Design
In some Non-LLMs, such as 6DRepNet and HopeNet, cropping the human head region is required as the initial step. In this paper, a new prompt method is proposed, allowing us to train and predict HPE utilizing the information of full images. In our prompts, BBox coordinates are leveraged to specify the human head of interest when multiple people are present. Therefore, the system is capable of effectively focusing on specific heads, which makes it easier to reduce the need for labor-intensive manual annotations and automate the inference process.
Meanwhile, the global features from self-attention and head of interest features from cross-attention are both learnt to improve the robustness of HPE task. Figure 1 (b) shows an example of our designed prompts and responses for the HPE task. More prompt details are in Appendix A.1 Table. 5.
4.2 Datasets
Table 1: A detailed overview of various datasets used in our framework.
Task Dataset # of Images # of Heads Usage
Train Test Train Test
Weak Label Images CrowdHuman shao2018crowdhuman 11,731 - 94,795 - Stage 1
Task-specific Images Agora patel2021agora 9,654 - 64,187 - Stage 2, 4
Rehearsal Images Refcoco yu2016modeling 42,404 3785 - - Stage 4, 5
Refcoco+ yu2016modeling 42,278 3773 - - Stage 4, 5
Refcocog mao2016generation 42,226 5023 - - Stage 4, 5
Evaluation Images CMU Panoptic joo2016panoptic - 16,216 - 32,738 Stage 5
Table 1 outlines datasets used in various stages in our framework. The CrowdHuman dataset shao2018crowdhuman serves as the pre-training dataset due to its extensive collection of human images. Its head pose annotations are derived from pseudo-labels inferred by the pre-trained 6DRepNet Hempel_2022 model, which is referred as the weak label images. The synthetic Agora dataset patel2021agora serves as the fine-tuning HPE dataset, which encompasses full-range of human head yaw angle images and provides the GT of SMPL-X parameters pavlakos2019expressive . Its head pose annotations are generated using the method of DirectMHP zhou2023directmhp . The Refcoco yu2016modeling , Refcoco+ yu2016modeling , and Refcocog mao2016generation train datasets, which are originally utilized by CogVLM to learn BBox prediction, are chosen as rehearsal images to help mitigate the catastrophic forgetting of existing knowledge. In our experiments, various portions of the rehearsal images are applied to determine the optimal rehearsal ratio scialom2022finetuned ; huang2024mitigating to address the catastrophic forgetting problem. A subset of the CMU Panoptic dataset (CMU dataset) is selected as the test dataset for evaluating HPE task, as its panoptic images of real people closely mirror real-life scenarios. The selection of images and annotations is guided by the DirectMHP  [^2] https://github.com/hnuzhy/DirectMHP. To evaluate object BBox localization, the test datasets including testA and testB data from Refcoco and Refcoco+, as well as the test dataset from Refcocog, are selected as the BBox evaluation datasets.
4.3 Implementation Details
[…]
4.4 Evaluation Metrics
Angel Error Ratio ($E_{angle}$): $E_{angle}$ = $\frac{e_{angle}}{t_{angle}}$, where $e_{angle}$ denotes the number of invalid HPE answers and $t_{angle}$ denotes the number of total HPE answers. This new metric is defined to assess the capability of models to provide relevant numerical outputs for HPE task. When we prompt with a HPE query, the CogVLM could produce irrelevant responses such as an natural language processing (NLP) task response like “a person head”, a bounding box task response like “[[111,222,333,444]]”, or even a completely nonsensical response like “111,999,999,999…”.
BBox Error Ratio ($E_{bbox}$): $E_{bbox}$ = $\frac{e_{bbox}}{t_{bbox}}$, where $e_{bbox}$ denotes the number of invalid BBox answers and $t_{bbox}$ denotes the number of total BBox answers. This new metric is defined to assess the capability of models to provide relevant numerical outputs for BBox prediction task.
BBox accuracy (ACC.): Acc. = $\frac{m}{\hat{m}}$, where $m$ denotes the number of valid BBox answers with IoU > 0.5 and $\hat{m}$ denotes the number of total valid BBox answers. A BBox prediction is considered to be accurate if the intersection over union (IoU) between the GT and the prediction exceeds 0.5 yu2016modeling . And the invalid answers are excluded from accuracy and MAE calculation.
MAE of Euler angles (MAE): For HPE task, the MAE of GT Euler angles and predicted Euler angles is defined as follows:
where $\hat{A}_{i}$ represents the GT’s Euler angles, $A_{i}$ represents the predicted Euler angles, and variable $n$ denotes the number of valid HPE answers. The MAE in HPE task is different from its conventional definition, where the error is measured in a circular manner rather than linearly, leading to the inclusion of a term that minimizes the difference between the predicted and actual angle by considering a full 360-degree rotation Hempel_2022 ; zhou2023directmhp . In this paper, the MAE value is considered as the average of MAE for yaw, pitch and roll Euler angles.
4.5 Baseline Methods
Non-LLMs, including 6DRepNet, HopeNet and WHENet are selected as HPE Non-LLM baselines. The 6DRepNet model, recognized as the current SOTA, is specifically retrained and tested on the same Agora and CMU datasets used in the LLM experiment to ensure a fair comparison. This model is trained 100 epochs, and the best MAE is selected for baseline analysis with our HPE-CogVLM. The pre-trained models of HopeNet and WHENet are utilized because HopeNet scripts are hard-coded and WHENet training scripts are not publicly available.
Non-merging CogVLM, direct fine-tuned model without applying model merging technique, is selected to compare the effectiveness of our merging approach with fine-tuning only method scialom2022finetuned ; huang2024mitigating ; luo2024empirical . The difference of Non-merging CogVLM and our HPE-CogVLM methods is that the Non-merging CogVLM bypasses stages 2 and 3, instead it undergoes significantly more training iterations in stage 4 which is equal to the total iterations of stages 2 and 4 in the HPE-CogVLM framework. For examples, our HPE-CogVLM is fine-tuned 25k and 5k iterations on stage 2 and 4 respectively, while the Non-merging CogVLM is solely fine-tuned on stage 4 for 30k iterations. This ensures fair comparison with respect to HPE task training iterations.
Task Arithmetic (TA) merging CogVLM, which adheres to our framework but replacing the layer-based merging with TA based merging, is to provide a baseline for comparing our merging approach with another merging method. The TA merging process is chosen as it forms the foundation for many other merging algorithms yadav2024ties ; yu2023language . In this process, we set the lambda parameter of task arithmetic to 0.5 ilharco2023editing , assigning equal importance to both the BBox prediction task and the HPE task.
5 Experimental Results
5.1 Baseline Comparison
Table 2: Comparison of HPE-CogVLM performance with various baselines. The best results for each model are highlighted in bold.
Model Refcoco Refcoco+ Refcocog CMU Panoptic
$\boldsymbol{Acc_{test}}$ $\boldsymbol{E_{bbox}}$ $\boldsymbol{Acc_{test}}$ $\boldsymbol{E_{bbox}}$ $\boldsymbol{Acc_{test}}$ $\boldsymbol{E_{bbox}}$ $\boldsymbol{MAE_{test}}$ $\boldsymbol{E_{angle}}$
WHENet - - - - - - 29.55 -
HopeNet - - - - - - 22.16 -
6DRepNet - - - - - - 10.74 -
Original Grounding CogVLM 91.4% 0% 86.7% 0% 90.2% 0% - -
Non-merging CogVLM 91.1% 0% 85.2% 0% 88.9% 0% 8.18 0.13%
TA merging CogVLM 89.5% 0% 82.3% 0% 86.1% 0% 7.72 68.9%
HPE-CogVLM 90.5% 0% 84.7% 0% 87.8% 0% 7.36 0.052%
The results in Table 2 show the performance comparison between our HPE-CogVLM model and various baselines described in Section 4.5. In comparison with Non-LLMs, our HPE-CogVLM presents significantly lower MAE. The HPE-CogVLM MAE is 75.1%, 66.8%, and 31.5% lower than WHENet, HopeNet, and 6DRepNet, respectively. The Non-LLMs also perform worse than other CogVLM-based models, which highlights the superior robustness of VLM-based models over Non-LLMs.
In comparison with Non-merging CogVLM, HPE-CogVLM MAE is 10% lower than the Non-merging CogVLM. Meanwhile the $E_{angle}$ of our model is 2.5 times smaller than the Non-merging CogVLM. This indicates that our layer-based merging method is more proficient in HPE than the method that does not utilize any model merging technique. Regarding the BBox results, the HPE-CogVLM’s BBox prediction accuracy in test datasets is 0.6%, 0.5% and 1.1% lower than the Non-merging CogVLM, however, the Non-merging CogVLM costs five times more iterations for training rehearsal dataset as discussed in Section 4.5.
Comparing with TA merging CogVLM, our HPE-CogVLM wins in all metrics. For instance, when evaluated on test datasets, the BBox prediction accuracy of the HPE-CogVLM exceeds that of the TA merging CogVLM by 1%, 2.4%, and 1.7%, respectively. And the $E_{angle}$ of TA merging CogVLM is 68.9%, which is 1325 times greater than that of HPE-CogVLM, indicating that only 31% of the responses for the HPE task are valid. Due to the high number of invalid HPE responses, the MAE metric becomes ineffective for assessing the performance. This highlights that even with an additional round of fine-tuning, the task arithmetic merging fails to produce relevant numerical responses within our research domain, ultimately proving ineffective for the HPE task.
5.2 Performance of HPE-oriented CogVLM on HPE Task Only
Table 3: The HPE-oriented CogVLM model exhibits the highest HPE performance within our framework. The best results for each model are highlighted in bold.
Model Epochs $\boldsymbol{Acc_{test}}$ $\boldsymbol{MAE_{test}}$ $\boldsymbol{MAE_{training}}$ $\boldsymbol{E_{angle}}$
6DRepNet 3 - 12.70 9.40 -
6DRepNet 6 - 12.76 8.80 -
6DRepNet 9 - 11.44 7.90 -
6DRepNet 50 - 11.37 2.91 -
6DRepNet 100 - 11.4 2.23 -
HPE-oriented CogVLM 3 8.8% 6.40 - 0.0092%
HPE-oriented CogVLM 6 12.6% 6.31 - 0%
HPE-oriented CogVLM 9 11.0% 6.24 - 0%
In our framework, the HPE-oriented CogVLM is our most effective model for the HPE task only. Table 3 presents comparative performance results of 6DRepNet and the HPE-oriented CogVLM, both not accommodate BBox prediction capabilities, over similar training epochs. The low Refcoco test accuracy of HPE-oriented CogVLM is expected, given that no data rehearsal is implemented in this stage. In terms of MAE metric, the HPE-oriented CogVLM displays a gradual decrease in MAE from 6.4 at 3 epochs to 6.24 at 9 epochs. When compared our model with 6DRepNet, in the same epoch, our MAE shows much lower numbers than 6DRepNet. For example, in epoch 9, MAE of HPE-oriented CogVLM is 6.24 which is 45.5% lower than 6DRepNet. After extending the 6DRepNet training to 100 epochs, while its training MAE decreases from 9.40 to 2.23, the MAE on the CMU dataset does not improve, remaining stable around 11.4. This indicates that the model is over-fitting to the Agora dataset, with no enhancement in cross-dataset inference performance. This stark difference emphasizes the superior performance of VLM than Non-LLMs.
5.3 Selecting Optimal Rehearsal Ratios for Mitigating the Catastrophic Forgetting Problem
Table 4: Performance of weak label CogVLM under various rehearsal ratios.
Iterations Rehearsal Ratio $\boldsymbol{Acc_{test}}$ $\boldsymbol{E_{bbox}}$ $\boldsymbol{MAE_{test}}$ $\boldsymbol{E_{angle}}$
0k 0% 91.4% 0% - -
10k 0% 21.8% 0.026% 17.20 0.48%
10k 1% 77.5% 0.19% 21.51 0.85%
10k 10% 91.0% 0% 19.32 0.32%
10k 25% 91.5% 0% 19.92 0.23%
Table 4 presents the performance of weak label CogVLM across different proportions (0%, 1%, 10%, 25%) scialom2022finetuned ; huang2024mitigating ; luo2024empirical of the rehearsal dataset. The primary aim is to determine the appropriate data rehearsal ratio to retain old knowledge for the fine-tuning in stage 4. The Refcoco test accuracy at iteration 0 is 91.4%, indicating proficiency with the BBox prediction tasks. After the training is finished, the results demonstrate a clear trend that as the rehearsal ratio increases, the Refcoco test accuracy substantially improves. Starting at a low of 21.8% when no Refcoco data is used, the accuracy spikes to 77.5% with just 1% of rehearsal ratio, eventually reaching over 91% with 10% and 25% of rehearsal ratio. This clearly shows that the more original task data used in learning a new task, the less catastrophic forgetting occurs. In the $E_{bbox}$ column, the consistently low $E_{bbox}$ values suggest that the availability of BBox predictions tend to stabilize after 10K iterations. MAE and $E_{angle}$ for HPE task show a fluctuating trend. Since the head pose pseudo-label is provided for this pre-training experiment, they may not serve as a reliable metric for analysis. Rehearsal ratios of 10% and 25% are selected for the stage 4 experiment due to high Refcoco test accuracy. These numbers are significantly higher than the commonly used 1% rehearsal ratio in non-grounding tasks.
5.4 The Influence of Rehearsal Ratios on Multi-task Learning
(a) The influence of rehearsal ratio on MAE.
(b) The influence of rehearsal ratio on BBox Prediction.
Figure 3: (a) shows the MAE results under rehearsal ratio 10% and 25% on various models. (b) shows the Refcoco Test BBox accuracy results under rehearsal ratio 10% and 25% on various models.
Figure 3 presents comparative results for both the BBox prediction task and the HPE task under different rehearsal ratios. Between the two HPE-CogVLM models, the one with a lower rehearsal ratio (10%) achieves an MAE of 7.36, which is 12% lower than the 8.36 observed with the higher (25%) ratio. Conversely, the Refcoco test accuracy improves slightly with higher rehearsal ratios, showing increases of 0.3% compared to the lower ratio. The similar phenomenon also presents in Non-merging CogVLM and TA merging CogVLM results. Intuitively, a higher rehearsal ratio helps retain the existing knowledge better since more data from previous tasks is included in the fine-tuning process. Therefore, a higher rehearsal ratio improves existing knowledge retention but at the expense of new task performance. So we seek for balance between the retention of old knowledge and the performance on new tasks. In our case, the HPE-CogVLM with a 10% rehearsal ratio clearly stands out as the best model in both HPE and BBox prediction. More experiment results are cataloged in Appendix A.4 Table 7.
6 Conclusions
In our paper, we propose a novel framework to enhance the HPE task by using the grounding CogVLM. We design prompts to enable CogVLM to learn HPE from full images, explore the optimal rehearsal ratio to prevent catastrophic forgetting problem and introduce a layer-based merging method. This new framework exhibits outstanding robustness and effectiveness over both Non-LLMs and other VLMs based methods.
Limitations This paper is limited by GPU resources, which constrains the extent of our experiments. As a result, larger-scale experiments could not be fully explored within the scope of this study.
Appendix A Supplemental Material
A.1 Details of the Designed Prompts
Table 5: Prompts and Responses Design for HPE Task.

### v2

@@ III-A Stage 1: Pre-training of the Original Grounding CogVLM on Weak Label Data
To achieve this, the original grounding CogVLM undergoes pre-training on the CrowdHuman dataset, which consists of images of real people in diverse scenes. Since the original CrowdHuman dataset provides only accurate head BBox annotations and lacks ground truth (GT) annotations for HPE, we infer the weak HPE annotations using 6DRepNet. Therefore, the output model from this stage is termed as weak label CogVLM as shown in Figure 2. As the CrowdHuman dataset offers a rich collection of human head images in various poses and contexts, this approach enables the model to accurately locate the BBoxes of human heads and establishes a foundational understanding for the subsequent HPE task.
@@
III-B Stage 2: Supervised Fine-tuning of the Weak Label CogVLM on Task-specific (HPE) Data
Following the pre-training stage, the model progresses to a supervised fine-tuning phase which is exclusively trained on task-specific HPE images. As shown in Table I, we utilize the Agora dataset as the task-specific images. Unlike the broader CrowdHuman dataset used previously, the agora dataset consists of synthetic images that offer more accurate annotations tailored to capture precise head pose information. The detailed annotations in this dataset allow the model to focus on subtle variations in head orientation, enhancing its ability to make fine-grained predictions. This stage focuses on improving the HPE accuracy, addressing the weaknesses in the weak label model’s performance due to lower-quality annotations. The output model is referred as the HPE-oriented CogVLM as shown in Figure 2. This refined model combines the real-image exposure and contextual background knowledge gained from the previous stage with the task-specific precision acquired in this stage.
@@ III-B Stage 2: Supervised Fine-tuning of the Weak Label CogVLM on Task-specific (HPE) Data
III-C Stage 3: Layer-based Merging between Original Grounding CogVLM and HPE-oriented CogVLM
@@
During this key stage, the original grounding CogVLM is merged with HPE-oriented CogVLM (from Stage 2) based on cosine similarity criteria. In our framework, cosine similarity is calculated as the average cosine similarity between the layer parameter tensors of the original grounding CogVLM and the HPE-oriented CogVLM along the last dimension. Cosine similarity is used to gauge the amount of information shared between layers. The threshold of cosine similarity can help to determine whether layers from the HPE-oriented CogVLM should be integrated into the final model. Since LoRA fine-tuning is applied in previous stages, most of the original model parameters are only minimally altered, necessitating a high threshold for cosine similarity. In our experiments, We empirically set the threshold at 0.95. If the similarity falls below this threshold, we opt to completely retain the original knowledge. Otherwise, if the similarity exceeds the threshold, which indicates a substantial overlap in information due to the stringent criteria, we select the entire layer from the HPE-oriented CogVLM to guarantee the minimal risk of losing important existing knowledge. Consequently, the merging criteria is detailed as below:
@@
  • When the cosine similarity between two layers from each model is less than the threshold, we also select the layer from the original grounding CogVLM.
@@
Precedent methods usually involve in directly LoRA fine-tuning or merging models at the parameter-level by setting hyper parameters and developing algorithms to discard and merge specific parameters [47, 51, 52, 53]. However, we empirically find that directly LoRA fine-tuning does not achieve the desired HPE accuracy, and existing merging methods can achieve better accuracy but struggle when handling multiple grounding tasks with different output formats. They often blend output structures, which leads to invalid answers. For example, when we query with HPE prompts, the merged model may return a NLP response like “a person of head” or provide nonsensical responses like “[[999,231,123,389}”. More examples are detailed in Table II. To address these issues, we set a high cosine similarity threshold and adopt a ”winner-takes-all” method to select entire layers from either the original grounding CogVLM or the HPE-oriented CogVLM. This enables the merged model’s attention to align with the new HPE task at minimal cost. By applying a high similarity threshold, we introduce new knowledge only when there is substantial informational overlap between the two models, ensuring that even when a layer is chosen from the HPE-oriented model, existing knowledge remains well-preserved. This approach minimizes attention loss while also aligning the model effectively with the new task. Although this new capability introduced at this stage may be basic, they lay a strong foundation for further refinement, allowing the model to build on this baseline effectively in stage 4. Additionally, by incorporating entire layers rather than individual parameters, the method preserves the integrity of the model’s structure, reducing the risk of response blending when handling multiple grounding tasks with varying requirements.
III-D Stage 4: Continual Fine-tuning of Layer-based Merging CogVLM on Mixture Data
After merging, the layer-based merging CogVLM undergoes an additional round of fine-tuning with both the task-specific HPE images and the rehearsal images. The optimal rehearsal ratio for training rehearsal images is pre-defined during Stage 1 by running several parallel experiments. In these experiments, we tune the original grounding CogVLM with weak label images, each combined with varying proportions of rehearsal images (0%,1%,10%,25%). The rehearsal ratio that yields the best performance is then used to fine-tune the merged model in this stage. Unlike the fine-tuning in stage 2, this phase involves only a brief period of fine-tuning, less than one epoch. The rationale for incorporating additional brief fine-tuning is that while layer merging maintains parameter integrity and optimize model’s focus, it lacks the fine-tuned parameters necessary to enhance HPE prediction accuracy. In our approach, the merging model can be quickly fine-tuned to deliver accurate numerical predictions. The final output model of this stage is HPE-CogVLM as shown in Figure 2.
III-E Stage 5: Evaluation of HPE-CogVLM on Test Data
@@
Compared to directly LoRA fine-tuning and some model merging methods, this framework ensures that HPE-CogVLM maintains a high level of accuracy in HPE task, effectively mitigating the issues of blended invalid outputs and catastrophic forgetting problem. Experimental results will be shown in the section V.
IV Experiments Setup
IV-A HPE Task Prompt Design
TABLE II: Prompts and Responses Design for HPE Task.